Find the eccentric angle of the point (x1,y1) on the ellipse x2/6 + y2/2=1 whose distance from the centre is 2.

Grade:11

3 Answers

Rahul Thakur
7 years ago
let the eccentric angle be t ;   a= 6  ;  b2 = 2
x1= acost     x2= asint
(acost)+ (bsint)2  =4
 
t  = +45 ; −45
Shubhranshu Pattnaik
6 years ago
Let t be the eccentric angle , a2=6 , b2=2 and point p(acost,asint).
It’s given distance of point from center(0,0) is 2 so by using distance formula, 
                          (acost)2+(bsint)2=4 
                           On solving we get :
                          cos2t=1/2 
                         So t= 45
ankit singh
3 years ago
We know, 6Π+θ=π
θ=65π in anticlockwise direction
So, eccentric angle is 65π
thank u regards
ankit singh manit bhopl